# Fundamental mechanisms of high-electric-field vacuum arc initiation

### v1

@@ III.8 Post-test microscopy
All signals described in the previous sections are measured while the systems are exposed to high fields . However, a common feature of all these signals is that they offer indirect observations regarding their sources. Various attempts to directly observe variations in features and properties of metal samples have been performed, with the objective of trying to identify features that give a field enhancement fact and pertaining to the processes leading up to breakdown [6]. Preliminary attempts at measuring field-emitted electron currents inside a scanning electron microscope (SEM) and observing the emitting sites prior to breakdown did not show any significant surface feature leading to the breakdown, at the resolution level permitted by the SEM [121, 120].
@@
Such dependencies can be reproduced in simulations of late-stage plasma evolution and theoretical results in that respect reproduced experimental observations of damage properties as described in section VII. An example of the surface damage is provided in Fig. 36 where crater and subsurface large-scale plastic evolution are demonstrated. It is clear from the cut below the surface that the energetic process that led to the formation of the surface crater also caused plastic deformation, seen as grain refinement, below the surface at dimensions similar to the radius of the crater. This damage was observed in a dedicated dc experiment with large energies available for plasma discharge. However, the observed damage in radio frequency systems is limited by the amount of power delivered to the breakdown plasma. However, these craters are the result of the energy deposition during plasma discharge and, as such, do not hold direct evidence of processes leading to a breakdown onset. Moreover, they may serve to erase any such feature responsible that can be associated with pre-breakdown signals.
Typically, the surfaces of radio-frequency structures and direct-current electrodes are scrutinized post-mortem using microscopic techniques after the were exposed to a number of breakdowns to address questions related to the nature of conditioning (both in terms of surface and material states) as well as to the insights into the plasma composition. The most prominent features identified on surfaces that underwent breakdowns are craters formed by deposited material from the hot plasma and material displaced due to high temperature and pressure exerted by plasma on the surface.
@@ IV Plastic material response
Even though vacuum breakdown and arcing are dominated by plasma dynamics, the latter, in turn, is a consequence of other processes developing after plasma onset. Therefore, we may have various processes at the origin of a breakdown with similar late-stage plasma evolution characteristics. Here we divide breakdowns into three different groups according to their origin: extrinsic, intrinsic, and secondary events. Extrinsic events are those caused by external sources such as contaminates and singular pre-existing surface irregularities that may cause localized sudden bursts of currents under the effect of an applied field [59, 93, 94]. However, such extrinsic sources are known to be removed efficiently during surface conditioning and are, therefore, ruled out as controlling the long-term asymptotic dynamics, as the breakdown rates tend to stabilize at a specific rate. Furthermore, attempts to correlate observations of such features using microscopy before field exposure with the locations of breakdown craters following exposure were unsuccessful. It is especially noteworthy that the location of breakdown craters on large-grained OFE copper do not correlate with the most pronounced topographical features, such as grain boundaries and triple junctions.
Another differentiation between observed events is the distinction between fast versus slow dynamics; see section III.3. These two distinct time distributions, observed in dark currents and breakdowns, were suggested to correspond to two different processes. The slow dynamics represent the underlying intrinsic breakdown processes, while the fast dynamics result from the secondary follow-up events, i.e. events that are likely triggered by remnants of the preceding breakdown event [81, 150]. According to this hypothesis plastic evolution within the surface mainly affects initial intrinsic events, so-called primary breakdowns. Thus, observations and models related to these do not apply to breakdowns caused by extrinsic mechanisms, i.e. by contaminates or external damage, even if due to a follow-up to a previous event.
@@ IV.1 Link of breakdown rates to material response
Previous chapters established that the properties of metals, specifically those used as cathodes in DC systems, control the ability of the metal to withstand high electric fields (see Fig. 28)[32, 35]. It was demonstrated that different metals start developing conducting plasmas above their surfaces in fields much higher than those of the others. Although it was possible to rank different materials in order of the ascending breakdown fields (the average minimum electric fields resulted in an immediate breakdown for a given material), the mechanism by which this effect takes hold remained unexplained. Surprisingly, the thermodynamic properties such as melting temperature and elastic constants failed to demonstrate the observed correlation [23].
@@
Following the initial proposal in Ref. [127] for a mechanism related to defect nucleation, led to a further search for mechanisms based on sub-surface plastic evolution in materials exposed to high electric fields and leading to the full plasma onset. One realization of such a process can be the motion of pre-existing and multiplying extended defects - namely dislocations. This proposal corresponds well with the structural correlation observed in Fig. 28 [32, 35] as dislocation mobility is strongly affected by the metal crystalline structure through its effect on properties such as the number of easy glide planes, the low-energy dislocation structure, dislocation mobility, and the mechanism for release of sessile systems.
@@ IV.2 Plastic processes due to dislocations - from quasi-static to dynamic evolution
Dislocation dynamics is known to control plastic deformation of metals [9]. This dynamics is stochastic and the deformation depends on the finite probability of releasing various dislocation structures; thus, plastic deformation can occur below the deterministic yield point of the metal [125] with a diminishing probability as driving forces are decreased. The stochastic nature of the process is significantly pronounced when the low stresses are applied as a cyclic load, which offers the repetitive sampling of the finite probability of dislocation release. The collective dynamics of dislocations, which are highly affected by interactions between mobile and sessile ones, can lead to strain localization along specific slip systems [105]. These are known as persistent slip bands and can create significant protruding structures on the surface at the intersection between the slip band and the surface [107]. Such micron-scale features were repeatedly observed in fatigued copper structures using SEM [3].
@@
The structures identified in these cases involve stable extended periodic structures that can be observed after formation using SEM [13, 142]. However, detailed studies of surfaces exposed to high electric fields do not show similar features. Even conditioned surfaces (i.e., exposed to a long repetitive set of high-field pulses as described in Sec. III.4) retained their as-produced original global surface structure as exemplified in Fig. 38, This is markedly different from the surfaces where persistent slip bands evolved, as shown in Fig. 11 in [142]. In addition, if a breakdown indeed nucleated due to gradual deterioration of the surface or accumulation of damage leading to the growth of surface features, it should have been possible to identify the marking of these features outside the specific region of the breakdown itself. The gradual evolution over a wide area should produce independent sub-failure features also in similarly field-effected areas that are outside the specific breakdown nucleation site. As such features are not observed (see Section III.8), a causal relationship between plastic surface damage and strong variation in surface response to the electric field is expected to be linked to a sudden localized event. That such features are not observed by post-exposure microscopy is consistent with measured dark currents from the surface which are indicative of a stable surface structure. While field exposure does affect surface resilience to breakdown (the aforementioned conditioning effect) gradual accumulation of damage leading to breakdown, even if unobserved using post-exposure microscopy, would lead to a strong memory effect in breakdown rates even in conditioned surfaces, as the accumulated damage reduces the amount of additional damage needed for future breakdown nucleation. Such surface deterioration in conditioned surfaces is not observed and in line with all of the above, it is assumed that gradual evolution can not be used to describe intrinsic breakdown nucleation.
@@
On the other hand, these observations are consistent with an abrupt breakdown nucleation process. Such abrupt processes related to plastic response were described by a critical transition in the controlling parameter. As most experimental work was done on OFE Cu where plasticity is controlled by the highly mobile dislocations a theory based on collective motion of these defects was offered to explain the mechanism leading to the observed critical transition.
@@
Various efforts were made to explore the relation between applied fields and dislocation activity, which depends on the activation of preexisting sources, commonly sessile dislocation arrangements. This link may be due to the activation of the dislocation source by temperature, external stress, and variations in internal stresses. A significant effort to elucidate the mechanisms of dislocation dynamics within the copper surface exposed to a high electric field was made in a number of studies using atomistic simulations coupled with the electrostatic field effects [139, 141, 140, 176, 177, 11]. The incorporation of the latter in a fully consistent model describing atom-atom interactions allowed for valuable insights on possible mechanisms of dislocation dynamics within the metal surface under external electric fields. For example, high tensile stresses exerted by the electric field on the surface were shown to lead to a deterministic formation of surface asperities by dislocation motion to the surface and that new dislocation sources at preexisting defects can be activated providing the self-reinforced growth mechanism for surface asperities [139, 141, 140]. However, the estimation of the stresses required in these small-scale simulations to trigger the dislocation nucleation processes remained significantly larger than the Maxwell stresses that a reasonable value of an applied electric field can exert on the exposed metal surface [188, 176, 177, 11] .
@@
It is worthwhile noting that even without forming a large-scale structure on top of the surface, dislocations moving below the surface may lead to significant modifications of electronic response to an externally applied field. Such effects were offered in the past as an explanation for the increase in corrosion rates of metal subjected to cyclic loading [98, 99]. However, as of today, a direct link between sub-threshold deterministic structure formation and possibly observable variation in field emission related to breakdown nucleation has not been demonstrated.
@@
At stresses lower than the threshold for deterministic immediate activation of plastic processes, stochastic activation of such processes occurs through non-linear interactions, which can lead to a critical transition in the density of dislocations. Such phenomena have been shown to control threshold plastic response in various scenarios [125, 53, 133] and even in brittle to ductile transitions [75]. The intermittent nature of plastic processes and the fact that such intermittent response can be connected with critical transition in dislocation mobility was reproduced in a mean-field model that uses the mobile dislocation density as the leading order parameter. [126, 148] The mobile dislocation density fluctuation (MDDF) model, is a mean-field kinetic model, which uses the number of moving mobile dislocations, and was constructed to reproduce the expected dynamics of driven surfaces under the stresses caused by an external field. [45] The model describes the kinetics of the creation and depletion of mobile dislocations in a single slip plane, neglecting interactions between slip planes and the spatial variation of the mobile dislocation density within one plane. In this model, mobile dislocations nucleate at existing sources, and their depletion is due to collisions with obstacles. The problem is formulated in terms of a birth-death master equation for the mobile dislocation population. The resulting model is unique in that, for the first time, it treats a breakdown in metals as a critical transition, due to the stochastic evolution of dislocations under an external field. In contrast with gradual (or linear) evolution models, the transition does not require the appearance of sub-breakdown surface features.
@@
It is noteworthy that such a model leads to the prediction that conditioning would involve evolution in the dislocation-related structures at the upper layer of the electrode. Indeed, microcopy of cross-sectional samples from conditioned electrodes, which were made from fine-grained hard OFE copper, showed the formation of dislocation-denuded zones with conditioning as demonstrated in figure 40[62].
@@
The MDDF leads to a non-trivial dependence of the mean time between breakdowns and the applied field . However, in two regimes, a simple approximated solution can be derived. For high fields, in the strongly driven state above the fluctuations controlled regime [45], the mean time between breakdowns is dominated by an exponent of leading to , as was previously suggested [127]. However, for lower fields, where breakdowns are fluctuation-driven, the full kinetic expression is approximated using a linear dependence of on ,
@@
[45]. These dependencies reproduce observed behavior in field and temperature dependencies of the measured saturation breakdown rate over a wide range of drive conditions, demonstrated in figure 39 [43, 45].
@@
In the described studies, the dislocations are modeled to respond to elastic fields in the material. The applied field modifies the stress generated in the material without a direct interaction between the electrons and the dislocation within the material. This assumption relies on studies that showed that only at very high current densities direct momentum transfer from the electrons to the dislocation provide significant contribution relative to the elastic interaction [159, 101]
@@ IV.4 Pre-breakdown fluctuations
A general characteristic of critical systems is that close the critical transition the system state fluctuates. As was observed in other systems where plastic behavior was controlled by a critical transition, fluctuations in plastic activity and specifically dislocation motion precede the critical transition [156, 53, 189]. In the case of pre-breakdown plastic activity direct monitoring of the plastic response within the exposed surface in-situ is still unavailable, and surface evolution is tractable only through in-direct measurements.
@@
In the MDDF model it was suggested that critical transition in the density of the mobile dislocations within a metal can lead to a rapid increase in local current [45, 43]. While this model is in line with previous suggestions that the electric field is enhanced at the breakdown nucleation site due to localized plastic activity [176, 11] it does not detail the specific of the plastic process. However, it can be speculated that independent of the specific mechanism, an increase in the pre-BD dislocations fluctuations, which are an essential part of the MDDF, would lead to an increase in field emission current. Therefore, although no observable plastic evolution of the surface needs to be observed, an increase in field emission fluctuations should be measured.
@@
In the model subcritical fluctuations of the mobile dislocation density might be observed as fluctuations of the dark current during routine operation [43, 45]. It is important to note that the rate of these fluctuations differs significantly from the breakdown rate and is not linearly related to it. Theoretical estimates of these pre-BD fluctuations were derived based on estimates of surface annihilation rates for mobile dislocations as predicted by the MDDF [44].
@@
For a study of these fluctuations, a measurement of the changes in the dark current, rather than its average value, were conducted in CERN and compared to theoretical estimates of these fluctuations as derived from the full stochastic description of the MDDF. Measurements were done in a dedicated setup and a sub-BD event identification served to identify the time distribution of the specific isolated dark-current spikes.
@@
Theoretical estimates of the dark current spikes were based on the assumption that these are correlated with multiplication events of the mobile dislocation population. Assuming, that all transitions within the model are individual Poissonian processes led to a theoretical description of the time distribution between current spikes which follows a two-parameter hypoexponential distribution with parameter values derived directly from the model [44]. the model reproduced well both distribution 41 and average fluctuations rate 42.
@@
Aside from providing support for the validity of an MDDF type model, an interesting prospect deriving from these observations is that this correlation may be utilized to reduce the breakdown rate in high-voltage setups and improve operational procedures in such systems, by using rates of observed dark current spikes to determine the level of conditioning of electrodes, or by monitoring spikes as observed in situ prior to breakdown.
@@
Aside from field emitted current spikes distribution, the fluctuations in the average field emitted currents were used in [87] to create a field-specific evaluation of the surface field enhancement factor . The results of this analysis are demonstrated in Fig 43, show insignificant variation of the field enhancement factor with the field. While these measurements are scarce and are yet to be performed up to the BD threshold itself [87], they do provide an additional indirect indication that field emission is not controlled by a gradual process, strengthening the hypothesis of a critical transition.

### v2

@@ III.8 Post-test microscopy
All signals described in the previous sections are measured while the systems are exposed to high fields . However, a common feature of all these signals is that they offer indirect observations regarding their sources. Various attempts to directly observe variations in features and properties of metal samples have been performed, with the objective of trying to identify features that give a field enhancement factor and pertaining to the processes leading up to breakdown [6]. Preliminary attempts at measuring field-emitted electron currents inside a scanning electron microscope (SEM) and observing the emitting sites prior to breakdown did not show any significant surface feature leading to the breakdown, at the resolution level permitted by the SEM [130, 129].
@@
Such dependencies can be reproduced in simulations of late-stage plasma evolution and theoretical results in that respect reproduced experimental observations of damage properties as described in Sec. VII. An example of the surface damage is provided in Fig. 37 and Fig. 38 where crater and subsurface large-scale plastic evolution are demonstrated. It is clear from the cut below the surface that the energetic process that led to the formation of the surface crater also caused plastic deformation, seen as grain refinement, below the surface at dimensions similar to the radius of the crater. This damage was observed in a dedicated dc experiment with large energies available for plasma discharge. The observed damage in radio frequency systems is instead limited by the amount of power delivered to the breakdown plasma. These craters are the result of the energy deposition during plasma discharge and, as such, do not hold direct evidence of processes leading to a breakdown onset. Moreover, they may serve to erase any such feature responsible that can be associated with pre-breakdown signals.
Typically, the surfaces of radio-frequency structures and direct-current electrodes are scrutinized post-mortem using microscopic techniques after they were exposed to a number of breakdowns to address questions related to the nature of conditioning (both in terms of surface and material states) as well as to insights into the plasma composition. The most prominent features identified on surfaces that underwent breakdowns are craters formed by deposited material from the hot plasma and material displaced due to high temperature and pressure exerted by plasma on the surface.
@@ IV Plastic material response
Even though vacuum breakdown and arcing are dominated by plasma dynamics, the latter, in turn, is a consequence of other processes developing after plasma onset. Therefore, we may have various processes at the origin of a breakdown with similar late-stage plasma evolution characteristics. Here, we divide breakdowns into three different groups according to their origin: extrinsic, intrinsic, and secondary events.
Extrinsic events are those caused by external sources such as contaminants and singular preexisting surface irregularities that may cause localized sudden bursts of currents under the effect of an applied field [62, 98, 99]. However, such extrinsic sources are known to be removed efficiently during surface conditioning and are, therefore, ruled out as controlling the long-term asymptotic dynamics, as the breakdown rates tend to stabilize at a specific rate.
Furthermore, attempts to correlate observations of such features using microscopy before field exposure with the locations of breakdown craters following exposure were unsuccessful. It is especially noteworthy that the location of breakdown craters on large-grained OFE copper does not correlate with the most pronounced topographical features, such as grain boundaries and triple junctions.
Another differentiation between observed events is the distinction between fast versus slow dynamics; see section III.3. These two distinct time distributions, observed in field emission currents and breakdowns, were suggested to correspond to two different processes. The slow dynamics represent the underlying intrinsic breakdown processes, while the fast dynamics result from the secondary follow-up events, i.e., events that are likely triggered by remnants of the preceding breakdown event [86, 159]. According to this hypothesis, plastic evolution within the surface mainly affects initial intrinsic events, so-called primary breakdowns. Thus, observations and models related to these do not apply to breakdowns caused by extrinsic mechanisms, i.e., by contaminants or external damage, even if due to a follow-up to a previous event.
@@ IV.1 Link of breakdown rates to material response
Previous chapters established that the properties of metals, specifically those used as cathodes in dc systems, control the ability of the metal to withstand high electric fields (see Fig. 29) [35, 38]. It was demonstrated that certain metals start developing conducting plasmas above their surfaces in fields much higher than those of the others. Although it was possible to rank different materials in order of the ascending breakdown fields (the average minimum electric fields resulting in an immediate breakdown for a given material), the mechanism by which this effect takes hold remained unexplained. Surprisingly, thermodynamic properties such as melting temperature and elastic constants failed to demonstrate the observed correlation [25].
@@
The initial proposal in Ref. [136] for a mechanism related to defect nucleation, led to a further search for mechanisms based on sub-surface plastic evolution in materials exposed to high electric fields and leading to the full plasma onset. One realization of such a process can be the motion of preexisting and multiplying extended defects - namely, dislocations. This proposal corresponds well with the structural correlation observed in Fig. 29 [35, 38] as dislocation mobility is strongly affected by the metal crystalline structure through its effect on properties such as the number of easy glide planes, the low-energy dislocation structure, dislocation mobility, and the mechanism for release of sessile systems.
@@ IV.2 Plastic processes due to dislocations - from quasi-static to dynamic evolution
Dislocation dynamics is known to control plastic deformation of metals [9]. These dynamics are stochastic and the deformations depend on the finite probability of releasing various dislocation structures. Thus plastic deformation can occur below the deterministic yield point of the metal [134] with a diminishing probability as driving forces are decreased. The stochastic nature of the process is significantly pronounced when the low stresses are applied as a cyclic load, which offers the repetitive sampling of the finite probability of dislocation release. The collective dynamics of dislocations, which are highly affected by interactions between mobile and sessile ones, can lead to strain localization along specific slip systems [112]. These are known as persistent slip bands and can create significant protruding structures on the surface at the intersection between the slip band and the surface [115]. Such micron-scale features were repeatedly observed in fatigued copper structures using SEM [3].
@@
The structures identified in these cases involve stable extended periodic structures that can be observed after formation using SEM [13, 151]. However, detailed studies of surfaces exposed to high electric fields do not show similar features. Even conditioned surfaces (i.e., exposed to a long repetitive set of high-field pulses as described in Sec. III.4) retained their as-produced original global surface structure as exemplified in Fig. 39. This is markedly different from the surfaces where persistent slip bands evolved, as shown in Fig. 11 in [151]. In addition, if a breakdown indeed nucleated due to the gradual deterioration of the surface or accumulation of damage leading to the growth of surface features, it should have been possible to identify the marking of these features outside the specific region of the breakdown itself. The gradual evolution over a wide area should produce independent sub-failure features also in similarly field-affected regions that are outside the particular breakdown nucleation site. As such features are not observed (see subsection III.8), a causal relationship between plastic surface damage and strong variation in surface response to the electric field is expected to be linked to a sudden localized event. That such features are not observed by post-exposure microscopy is consistent with measured field emission currents from the surface, which are indicative of a stable surface structure. While field exposure does affect surface resilience to breakdown (the aforementioned conditioning effect) gradual accumulation of damage leading to breakdown, even if unobserved using post-exposure microscopy, would lead to a strong memory effect in breakdown rates even in conditioned surfaces, as the accumulated damage reduces the amount of additional damage needed for future breakdown nucleation. Such surface deterioration in conditioned surfaces is not observed, and in line with all of the above, it is assumed that gradual evolution can not be used to describe intrinsic breakdown nucleation.
@@
On the other hand, these observations are consistent with an abrupt breakdown nucleation process. Such abrupt processes related to plastic response were described by a critical transition in the controlling parameter. As most experimental work was done on OFE Cu, where plasticity is controlled by the highly mobile dislocations, a theory based on the collective motion of these defects was offered to explain the mechanism leading to the observed critical transition.
@@
Various efforts were made to explore the relation between applied fields and dislocation activity, which depends on the activation of preexisting sources, commonly sessile dislocation arrangements. This link may be due to the activation of the dislocation source by temperature, external stress, and variations in internal stresses. A significant effort to elucidate the mechanisms of dislocation dynamics within the copper surface exposed to a high electric field was made in several studies using atomistic simulations coupled with the electrostatic field effects [148, 150, 149, 186, 187, 11]. The incorporation of the latter in a fully consistent model describing atom-atom interactions allowed for valuable insights on possible mechanisms of dislocation dynamics within the metal surface under external electric fields. For example, high tensile stresses exerted by the electric field on the surface were shown to lead to a deterministic formation of surface asperities by dislocation motion to the surface, and new dislocation sources at preexisting defects can be activated, providing the self-reinforced growth mechanism for surface asperities [148, 150, 149]. However, the estimation of the stresses required in these small-scale simulations to trigger the dislocation nucleation processes remained significantly larger than the Maxwell stresses that a reasonable value of an applied electric field can exert on the exposed metal surface [198, 186, 187, 11] .
@@
It is worthwhile noting that even without forming a large-scale structure on top of the surface, dislocations moving below the surface may lead to significant modifications of the electronic response to an externally applied field. Such effects were offered in the past as an explanation for the increase in corrosion rates of metal subjected to cyclic loading [105, 106]. However, as of today, a direct link between sub-threshold deterministic structure formation and possibly observable variation in field emission related to breakdown nucleation has not been demonstrated.
@@
At stresses lower than the threshold for deterministic immediate activation of plastic processes, stochastic activation of such processes occurs through non-linear interactions, which can lead to a critical transition in the density of dislocations. Such phenomena have been shown to control threshold plastic response in various scenarios [134, 56, 142] and even in brittle to ductile transitions [79]. The intermittent nature of plastic processes and the fact that such intermittent response can be connected with a critical transition in dislocation mobility was reproduced in a mean-field model that uses the mobile dislocation density as the leading order parameter. [135, 157] The mobile dislocation density fluctuation (MDDF) model, is a mean-field kinetic model, which uses the number of moving mobile dislocations, and was constructed to reproduce the expected dynamics of driven surfaces under the stresses caused by an external field. [48] The model describes the kinetics of the creation and depletion of mobile dislocations in a single slip plane, neglecting interactions between slip planes and the spatial variation of the mobile dislocation density within one plane. In this model, mobile dislocations nucleate at existing sources, and their depletion is due to collisions with obstacles. The problem is formulated in terms of a birth-death master equation for the mobile dislocation population. The resulting model is unique in that, for the first time, it treats a breakdown in metals as a critical transition, due to the stochastic evolution of dislocations under an external field. In contrast with gradual (or linear) evolution models, the transition does not require the appearance of sub-breakdown surface features.
@@
It is noteworthy that such a model predicts that conditioning would involve evolution in the dislocation-related structures at the upper layer of the electrode. Indeed, microscopy of cross-sectional samples from conditioned electrodes, which were made from fine-grained hard OFE copper, showed the formation of dislocation-denuded zones with conditioning as demonstrated in Fig. 40 [65].
@@
The MDDF leads to a non-trivial dependence of the mean time between breakdowns and the applied field . However, in two regimes, a simple approximated solution can be derived. For high fields, in the strongly driven state above the fluctuations controlled regime [48], the mean time between breakdowns is dominated by an exponent of leading to , as was previously suggested [136] where is a material and sample-dependent constant related to the plastic response of the surface to the applied field. However, for lower fields, where breakdowns are fluctuation-driven, the full kinetic expression is approximated using a linear dependence of on ,
@@
[48]. Note that for , the local field at the surface, the applied field is multiplied by a local field enhancement factor of the order of , represents the threshold for the specific process and is a pre-exponential factor depending on current emitting area These dependencies reproduce observed behavior in field and temperature dependencies of the measured saturation breakdown rate over a wide range of drive conditions, demonstrated in Fig. 41 [46, 48].
@@
In the described studies, the dislocations are modeled to respond to stress in the material. The applied field modifies the stress generated in the material without a direct interaction between the electrons and the dislocation within the material. This assumption relies on studies that showed that only at very high current densities, direct momentum transfer from the electrons to the dislocation provides a significant contribution relative to the elastic interaction [168, 108]
@@ IV.4 Pre-breakdown fluctuations
A general characteristic of critical systems is that close to the critical transition the system state fluctuates. As was observed in other systems where plastic behavior was controlled by a critical transition, fluctuations in plastic activity and specifically dislocation motion precede the critical transition [165, 56, 199].
In the case of pre-breakdown induced plastic activity, it is suggested that the system state is characterized by a surface property that affects the emitted pre-BD current and evolves through the motion of mobile dislocations. However, direct monitoring of the plastic response within the exposed surface in situ is still unavailable, and surface evolution is tractable only through indirect measurements.
@@
Along these lines, in the MDDF model it was suggested that a critical transition in the density of the mobile dislocations within a metal can lead to a rapid increase in local current [48, 46]. While this model is in line with previous suggestions that the electric field is enhanced at the breakdown nucleation site due to localized plastic activity [186, 11], it does not detail the specifics of the plastic process. However, it can be speculated that independent of the specific mechanism, an increase in the pre-BD dislocation fluctuations, which are an essential part of the MDDF, would lead to a rise in the field emission current. Therefore, although no observable plastic evolution of the surface needs to be observed, an increase in field emission fluctuations should be measured.
@@
In the model, subcritical fluctuations of the mobile dislocation density might be observed as fluctuations of the field emitted current during routine operation [46, 48]. It is important to note that the rate of these fluctuations differs significantly from the breakdown rate and is not linearly related to it. Theoretical estimates of these pre-BD fluctuations were derived based on estimates of surface annihilation rates for mobile dislocations as predicted by the MDDF [47].
@@
For a study of these fluctuations, measurements of the changes in the field emission current, rather than its average value, were conducted in CERN and compared to theoretical estimates of these fluctuations as derived from the full stochastic description of the MDDF. Measurements were done in a dedicated setup, and a sub-BD event identification served to identify the time distribution of the specific isolated dark-current spikes.
@@
Theoretical estimates of the dark current spikes were based on the assumption that these are correlated with multiplication events of the mobile dislocation population. The assumption that all transitions within the model are individual Poissonian processes led to a theoretical description of the time distribution between current spikes, which follows a two-parameter hypoexponential distribution with parameter values derived directly from the model [47]. The model reproduced well both the distribution and the average fluctuation rate, as demonstrated in Fig. 42 and in Fig. 43 correspondingly.
@@
Aside from providing support for the validity of an MDDF type model, an interesting prospect deriving from these observations is that this correlation may be utilized to reduce the breakdown rate in high-voltage setups and improve operational procedures in such systems, by using rates of observed dark current spikes to determine the level of conditioning of electrodes, or by monitoring spikes as observed in situ before breakdown.
@@
Aside from the field-emitted current spikes distribution, the fluctuations in the average field-emitted currents were used in [92] to create a field-specific evaluation of the surface field enhancement factor . The results of this analysis are presented in Fig. 44, demonstrating insignificant variation of the field enhancement factor with the field. While these measurements are scarce and are yet to be performed up to the BD threshold itself [92], they do provide an additional indirect indication that field emission is not controlled by a gradual process, strengthening the hypothesis of a critical transition.